### v1

# AR-KAN: Autoregressive-Weight-Enhanced Kolmogorov–Arnold Network for Time Series Forecasting

Abstract
[…]
Index Terms:
Time series forecasting, ARIMA, Kolmogorov-Arnold Network, KAN, Almost periodic functions
I Introduction
Time series forecasting is a fundamental task in signal processing[1][2], statistics[3], and numerous applied fields, including economics[4], meteorology[5], and healthcare[6]. Among classical approaches, the Autoregressive Integrated Moving Average (ARIMA) model[7] stands out as one of the most influential and widely adopted methods, because it integrates autoregression, differencing, and moving average elements to provide a comprehensible and effective approach for handling practical time series data, even when the time series is non-stationary.
Apart from the aforementioned statistics or Fourier analysis-based methods, neural networks have been utilized in time series forecasting for many years[8], with the goal of enabling the modeling of complex nonlinear dependencies. Architectures such as Multi-Layer Perceptrons (MLPs)[10] and Recurrent Neural Networks (RNNs)[11], particularly Long Short-Term Memory (LSTM) networks[12], have been widely studied. In recent years, Transformer-based models[13][14][15] have gained popularity due to their self-attention mechanism and parallel processing capabilities. Meanwhile, state space models like Mamba[16] have emerged as efficient alternatives to attention mechanisms, offering linear-time computation and strong performance on long-range sequences. More recently, Kolmogorov-Arnold Networks (KANs)[17][18] have been introduced as a novel architecture with high expressivity and flexible modeling of nonlinear mappings. In parallel, the rapid progress of large language models (LLMs) has led to approaches such as LLMTime[42] and Time-LLM[19], which adapt pretrained language models to temporal tasks by leveraging their strong generalization and sequence modeling capabilities.
In the context of neural forecasting, a specialized research focuses on spectral analysis through specific networks, such as Fourier Neural Networks (FNNs)[20]. These models incorporate Fourier series to enhance spectral modeling[21]. Representative examples include the Fourier Neural Operator (FNO)[23] and the Fourier Analysis Network (FAN)[22], which have been applied to physics-informed learning, partial differential equation solving, and time series prediction.
Nevertheless, these neural network models grounded in representation by Fourier series may overlook a key theoretical constraint: the additive combination of periodic elements does not necessarily result in a periodic function[24][25]. Throughout history, this important topic prompted N. Wiener to create the renowned Generalized Harmonic Analysis (GHA) theory, which works alongside the spectral analysis of time series. When the constituent frequencies are incommensurable, the resulting signal is almost-periodic[26], meaning that it exhibits recurrence without strict periodicity. Empirical studies show that for such signals, even advanced neural models, including FNNs, are often outperformed by classical ARIMA[27][28] and an evaluation could be referred as to our recent work [41].
Empirical studies indicate that, for such signals, even advanced neural models such as FNNs are often outperformed by classical ARIMA methods [27][28]. A more detailed evaluation can be found in our recent work [41].
To address this, we propose AR-KAN, a hybrid model that integrates the strengths of traditional and modern approaches. Based on the Universal Myopic Mapping Theorem[29][30], AR-KAN employs a KAN as the static nonlinear component, while introducing memory through a pre-trained autoregressive (AR) model. This design enables AR-KAN to combine the adaptability and expressiveness of KANs with the strong spectral bias inherent in traditional AR models. Furthermore, the AR memory module itself is a data-driven model whose weights are not fixed but are adaptively determined by the characteristics of the data. Additionally, it can be shown that when handling time series forecasting tasks, this module effectively eliminates redundancy while retaining the maximal amount of useful information. This property allows the model to flexibly adapt to various temporal patterns without manual intervention.
Our experiments confirm the effectiveness and generalizability of AR-KAN. On almost-periodic functions, it matches the performance of ARIMA. Moreover, across 18 real-world datasets from diverse domains, AR-KAN outperforms all baselines and achieves the best results on 72% of them (13 out of 18). These results highlight AR-KAN’s robustness, adaptability, and promise as a unified framework for time series forecasting.
The structure of this paper is organized as follows:
Section II introduces the background, including time series forecasting tasks, ARIMA, MLP, and KAN models. Section III presents the Universal Myopic Mapping Theorem and explains how it inspires the overall architecture of our AR-KAN model. Section IV describes the experiments conducted to demonstrate the effectiveness and generalizability of AR-KAN, including evaluations on two constructed almost-periodic functions and 18 real-world time series. Finally, Section V concludes the paper and discusses potential directions for future research.
II Background
II-A Time Series Forecasting and ARIMA
Time series forecasting aims to predict a sequence based on its past observations. Formally, given a univariate time series $\{x_{n}\}_{n=1}^{T}$, the forecasting problem involves learning a mapping $\mathcal{F}$ such that:
$\hat{x}_{n+h}=\mathcal{F}(x_{n},x_{n-1},\dots,x_{n-p+1}),$ (1)
where $\hat{x}_{n+h}$ denotes the forecast for $h$-steps ahead ($h=1$ in this paper), and $p$ is the order of historical dependence. This formulation can be extended to multivariate or probabilistic settings, but the central challenge remains: capturing the underlying temporal dynamics, dependencies, and possibly noise in the observed data.
A classical and widely used model for time series forecasting is ARIMA. ARIMA is particularly effective for stationary or differenced stationary processes. The general form of an ARIMA($p$, $d$, $q$) model is given by:
$\Phi(B)(1-B)^{d}x_{n}=\Theta(B)\epsilon_{n},$ (2)
where:
  • $B$ is the backshift operator, i.e., $B^{k}x_{n}=x_{n-k}$,
  • $\Phi(B)=1-a_{1}B-\dots-a_{p}B^{p}$ is the autoregressive (AR) polynomial of order $p$,
  • $\Theta(B)=1+b_{1}B+\dots+b_{q}B^{q}$ is the moving average (MA) polynomial of order $q$,
  • $d$ is the degree of differencing to ensure stationarity,
  • $\epsilon_{n}$ is assumed to be white noise: $\epsilon_{t}\sim\mathcal{N}(0,\sigma^{2})$.
The integration component $(1-B)^{d}$ transforms non-stationary series into stationary ones by differencing. The ARIMA model captures linear temporal dependencies and is known for its statistical interpretability and relatively low computational cost. Despite its simplicity, ARIMA remains a strong baseline in many practical applications, especially when the underlying signal exhibits regular, stationary behavior.
II-B MLP and KAN
MLP is one of the most fundamental architectures in neural networks. An MLP consists of multiple layers of affine transformations followed by pointwise nonlinear activations. Given an input $x\in\mathbb{R}^{d}$, an $L$-layer MLP computes:
$f_{\text{MLP}}(x)=W^{(L)}\sigma_{L-1}\cdots\sigma_{1}\left(W^{(1)}x+b^{(1)}\right)+b^{(L)},$ (3)
where $W^{(\ell)}$, $b^{(\ell)}$ are learnable parameters, and $\sigma_{\ell}$ denotes the nonlinear activation at layer $\ell$.
Figure 1: Universal Myopic Mapping Theorem.
Figure 2: Model Structure of AR-KAN.
However, MLPs exhibit a well-known spectral bias[31], meaning they tend to learn low-frequency components of the target function earlier and more accurately than high-frequency components. While this inductive bias can be beneficial in some applications, it limits the ability of MLPs to capture fine-grained or oscillatory patterns in data.
To overcome the limited expressiveness of fixed activation functions in traditional MLPs, KANs have been proposed as a more flexible and interpretable alternative. KANs are inspired by the Kolmogorov–Arnold representation theorem[32], which states that any multivariate continuous function $f:[0,1]^{d}\rightarrow\mathbb{R}$ can be expressed as a finite composition of univariate continuous functions:
$f(x_{1},\dots,x_{d})=\sum_{q=1}^{2d+1}\phi_{q}\left(\sum_{i=1}^{d}\psi_{qi}(x_{i})\right),$ (4)
where $\phi_{q}$ and $\psi_{qi}$ are univariate continuous functions. Inspired by this constructive result, KANs replace the fixed nonlinear activations in MLPs with learnable univariate functions, typically represented by splines.
Given an input $x\in\mathbb{R}^{d}$, an $L$-layer KAN computes:
$f_{\text{KAN}}(x)=\Phi^{(L)}\Psi^{(L-1)}\cdots\Psi^{(1)}(x),$ (5)
where each layer $\Psi^{(\ell)}:\mathbb{R}^{d_{\ell}}\rightarrow\mathbb{R}^{d_{\ell+1}}$ is defined by:
$\Psi^{(\ell)}(x)]_{j}=\sum_{i=1}^{d_{\ell}}w_{ij}^{(\ell)}\cdot\psi_{ij}^{(\ell)}(x_{i}),$ (6)
and $\Phi^{(L)}$ denotes the final output transformation, typically of the same form. Here, each $\psi_{ij}^{(\ell)}$ is a learnable univariate function, often implemented using splines, and $w_{ij}^{(\ell)}$ are learnable scalar weights.
Unlike MLPs, KANs do not exhibit a low-frequency spectral bias [33]. This enables them to capture high-frequency and oscillatory components more effectively, making them well suited for modeling time series with rich spectral structures.
However, this advantage can also introduce challenges. Without a low-frequency bias, KANs tend to be more sensitive to high-frequency noise[34] and may have difficulty learning functions with limited regularity[35]. In such cases, the model may overfit to spurious variations or become unstable during training.
Nevertheless, in most real-world time series, especially those with structured periodicity, seasonal trends, or non-stationary high-frequency patterns, this characteristic is beneficial. The ability of KANs to model a broad spectrum of frequency behaviors often leads to better performance compared to MLPs.
III AR-KAN
AR-KAN is derived from the Universal Myopic Mapping Theorem. Therefore, in this section, we first introduce the Universal Myopic Mapping Theorem, then followed by a detailed explanation of the AR-KAN model architecture.
III-A Universal Myopic Mapping Theorem
The Universal Myopic Mapping Theorem [29][30] provides a powerful theoretical guarantee for modeling dynamic systems using shallow, feedforward structures. Specifically, it states that any shift-invariant and myopic dynamical map can be uniformly approximated arbitrarily well by a two-stage architecture: a bank of linear filters followed by a static nonlinear mapping, as shown in Fig. 1.
Theorem 1 (Universal Myopic Mapping Theorem [29][30]).
Let $\mathcal{M}$ be a shift-invariant and myopic dynamical system that maps a real-valued time series $\{x_{n}\}_{n\in\mathbb{Z}}$ to outputs $\{y_{n}\}$ via a causal and bounded operator. Then, for any $\varepsilon>0$, there exists a finite collection of linear filters $\{h_{i}\}_{i=1}^{N}$ and a continuous static nonlinear function $f_{\theta}:\mathbb{R}^{N}\to\mathbb{R}$ such that the approximation
$y_{n}\approx f_{\theta}\left((h_{1}*x)_{n},(h_{2}*x)_{n},\dots,(h_{N}*x)_{n}\right)$
satisfies
$\sup_{n}\left|y_{n}-f_{\theta}\left((h_{1}*x)_{n},\dots,(h_{N}*x)_{n}\right)\right|<\varepsilon,$
where * denotes convolution and $(h_{i}*x)_{n}=\sum_{\tau}h_{i}(\tau)x_{n-\tau}$.
This theorem establishes that it is theoretically sufficient to model a wide class of dynamical systems using a finite bank of linear filters followed by a nonlinear function, without requiring recurrent or deep sequential architectures. The key property of myopia means that each output depends only on a bounded past history, and shift-invariance ensures time-homogeneity.
III-B Model Structure of AR-KAN
Inspired by the Universal Myopic Mapping Theorem, we design the AR-KAN as a two-stage architecture composed of a data-driven memory module and a static nonlinear mapping, as illustrated in Fig. 2. The static nonlinear network is implemented using a KAN, which has been discussed in Section II to possess stronger spectral modeling capabilities than traditional MLPs, particularly for high-frequency signals. For the memory module, we adopt a pre-trained AR model to serve as the bank of linear filters, effectively incorporating the strengths of classical linear time series models into our architecture.
The memory module operates in the following manner: we first train an AR model from the input time series $\{x(n)\}$ to predict the next step via
$\hat{x}(n+1)=\sum_{i=0}^{p-1}a_{i}x(n-i),$ (7)
where $p$ is the AR order and $\{a_{i}\}_{i=0}^{p-1}$ are the learned AR coefficients. These coefficients are then extracted to define a set of fixed linear filters. At each time step $n$, a delay buffer forms the historical input vector $\{x(n-i)\}_{i=0}^{p-1}$, which is multiplied elementwise with the corresponding $\{a_{i}\}_{i=0}^{p-1}$ and passed to the subsequent KAN module. This structure is equivalent to setting the impulse response of the $i$-th filter in Fig. 1 as:
$h_{i}(n)=a_{i}\,\delta(n-i),\quad 0\leq i\leq p-1,$ (8)
where $\delta(\cdot)$ is the Kronecker delta function.
To express the AR coefficients $\{a_{i}\}$ explicitly in terms of the time series $\{x(n)\}$, we can solve the Yule–Walker equations[36][37]. Specifically, let $\mathbf{a}=[a_{0},a_{1},\dots,a_{p-1}]^{\top}$ be the coefficient vector, $\mathbf{r}=[r(1),r(2),\dots,r(p)]^{\top}$ the autocorrelation vector, and $\mathbf{R}$ the $p\times p$ autocorrelation matrix given by
$\mathbf{R}=\begin{bmatrix}r(0)&r(1)&\cdots&r(p-1)\\ r(1)&r(0)&\cdots&r(p-2)\\ \vdots&\vdots&\ddots&\vdots\\ r(p-1)&r(p-2)&\cdots&r(0)\end{bmatrix}$, (9)
then the AR coefficients are computed via:
$\mathbf{a}=\mathbf{R}^{-1}\mathbf{r}.$ (10)
Here, the autocorrelation function $r(i)$ is defined as
$r(i)=\mathbb{E}[x(n)\,x(n-i)],$ (11)
or, in practice, estimated from the empirical data as
$r(i)\approx\frac{1}{N-i}\sum_{n=i}^{N-1}x(n)\,x(n-i),$ (12)
where $N$ is the total number of available samples.
This formulation reveals a key feature of our memory module: the filter weights $\{a_{i}\}$ are not fixed parameters, but are derived from the underlying data through statistical estimation. In contrast to static memory schemes such as tapped-delay lines[38] or gamma memory[39], our data-driven design allows the memory module to adapt flexibly to the autocorrelation structure of different time series.
III-C Analysis of the AR Memory Module
To further elucidate the advantage of the AR memory module, we provide a theoretical analysis demonstrating that it optimally preserves useful information while eliminating redundancy. Consider a general linear memory module with output:
$y_{i}(n)=w_{i}x(n-i),\quad 0\leq i\leq p-1,$ (13)
where $w_{i}$ are the weights.
We aim to maximize the total correlation between the memory outputs and the target $x(n+1)$, which represents the useful information captured:
$\text{max}\;\sum_{i=0}^{p-1}\mathbb{E}[y_{i}(n)x(n+1)].$ (14)
TABLE I: Test loss (MSE) of various models on Noisy Almost Periodic Functions
functions $\sigma$ ARIMA AR-KAN AR-MLP KAN MLP Transformer LSTM Mamba FAN FNO
$f_{1}$ 0.1 0.0142 0.0203 0.0270 0.1507 0.1216 0.0584 0.0743 0.1194 0.1173 0.0767
0.2 0.0550 0.0770 0.0959 0.1946 0.1273 0.3903 0.1462 0.2934 0.4266 0.1305
0.3 0.1206 0.1681 0.1999 0.2947 0.2408 0.4635 0.5209 0.3781 0.7023 0.1979
0.4 0.2155 0.2892 0.3543 0.6241 1.4625 1.5572 0.3932 0.5932 0.7965 0.7865
$f_{2}$ 0.1 0.0194 0.0193 0.0214 0.0515 0.1525 0.0947 0.0813 0.1149 0.0384 0.0322
0.2 0.0881 0.0724 0.0922 0.2812 0.1550 0.5346 0.2424 0.2593 0.5109 0.2747
0.3 0.1647 0.1593 0.1745 0.2577 0.6787 1.2197 0.4042 0.5592 0.3506 0.4277
0.4 0.3108 0.2769 0.3341 0.7100 1.1827 3.8209 0.4932 0.5914 0.7702 1.1133
Note: Bold numbers indicate the minimum value in each row; italic numbers indicate the second minimum value.
Figure 3: Performance of ARIMA, AR-KAN and FAN on Noisy Almost Periodic Functions, $\sigma=0.4$ (left: $f_{1}$, right: $f_{2}$).
However, this objective alone is insufficient, as it can be trivially maximized by arbitrarily increasing the magnitude of $w_{i}$, which would also amplify noise and irrelevant components. To prevent this and encourage the memory to focus on the most informative features, we introduce a constraint on the total output energy of the memory module:
$\text{min}\;\mathbb{E}\left[\left(\sum_{i=0}^{p-1}y_{i}(n)\right)^{2}\right].$ (15)
This constraint penalizes high-energy outputs, effectively forcing the memory to represent the target using a compact set of features and discard redundant information. We combine these two objectives into a single optimization goal:
$L=\sum_{i=0}^{p-1}\mathbb{E}[y_{i}(n)x(n+1)]-\frac{1}{2}\mathbb{E}\left[\left(\sum_{i=0}^{p-1}y_{i}(n)\right)^{2}\right].$ (16)
To find the optimal weights that maximize $L$, we solve $\frac{\partial L}{\partial\mathbf{w}}=0$ for $\mathbf{w}=[w_{0},w_{1},\dots,w_{p-1}]^{\top}$ gives:
$\mathbf{w}^{*}=\mathbf{R}^{-1}\mathbf{r},$ (17)
which is exactly the solution for the AR coefficients. This result confirms that the AR memory module optimally balances the dual goals of preserving predictive information and minimizing redundancy, providing a principled foundation for its use in AR-KAN.
This adaptability endows AR-KAN with stronger generalization across diverse temporal patterns. The linear filters capture data-specific short-term dynamics, while the nonlinear KAN component models higher-order, nonlinear interactions. Together, they form a powerful hybrid that balances interpretability, efficiency, and expressiveness in time series forecasting.
IV Experiments
We conduct experiments in two parts to demonstrate both the effectiveness and generalizability of AR-KAN. First, we perform experiments on noisy almost-periodic functions to show that modern models fall short of traditional ARIMA models in terms of spectral analysis, while our AR-KAN achieves performance comparable to ARIMA. Then, we extend the evaluation to 18 real-world datasets from Rdatasets[40], demonstrating that AR-KAN achieves the best performance on 72% of them. The detailed experimental settings are provided in appendices.
TABLE II: Test loss (MSE) of various models on Rdatasets
Datasets ARIMA AR-KAN AR-MLP KAN MLP LSTM FAN FNO LLMTime
a10_ts 0.1441 0.1353 0.4775 2.5033 2.2638 0.8809 0.4913 0.3851 0.3457
airpass_ts 0.3329 0.0706 0.0871 0.3046 0.3025 0.4249 0.5163 0.6982 0.1937
ausbeer_ts 0.0418 0.0357 0.0741 0.1031 0.5102 0.0692 0.0802 0.1114 0.0436
auscafe_ts 0.3301 0.3813 0.1180 2.6312 0.7564 0.3746 1.4820 3.4769 0.4463
BJsales_ts 0.3241 0.0032 0.0261 0.0358 0.7849 0.0643 0.2393 1.0370 0.0131
bricksq_ts 0.2080 0.0502 0.0823 0.2542 0.2769 0.2961 0.9625 0.2607 0.2541
co2_ts 0.0218 0.0014 0.0064 0.3079 0.0460 0.1640 0.1584 0.1963 0.0109
discoveries_ts 1.6030 2.1695 2.3091 1.7269 1.6264 1.7949 1.0153 0.8469 1.2922
economics_df_ts 3.5659 0.0845 0.4398 2.3047 1.8717 2.2520 1.8607 7.7670 1.6490
elec_ts 0.2731 0.0069 0.0060 0.093 0.0436 0.1833 0.3258 0.0625 0.0727
elecdaily_mts 0.4331 0.2123 0.2573 0.441 0.5366 0.6792 0.8981 0.5919 0.6127
elecequip_ts 0.3159 0.1528 0.1346 0.5968 0.5538 0.8984 0.7761 0.4870 1.4010
euretail_ts 0.4967 0.9964 1.3328 1.1984 0.4226 1.6740 1.1469 0.1821 1.5009
goog200_ts 4.7135 0.1228 0.8096 3.6632 3.5888 3.2584 3.0580 7.9012 1.1351
gtemp_both_ts 2.2374 0.2936 0.5328 3.2225 2.2946 1.6629 2.8678 1.6660 5.0938
h02_ts 0.2726 0.1263 0.1782 1.3708 0.5258 0.2103 0.6003 0.8209 0.1371
hsales2_ts 0.5781 0.5232 0.6301 2.1212 0.8286 1.7787 1.7065 0.8850 0.5667
hyndsight_ts 0.8729 0.2471 0.3961 1.5734 0.4892 0.5929 0.6164 0.6793 0.4510
Note: Bold numbers indicate the minimum value in each row.
Figure 4: Performance of ARIMA, AR-KAN and FNO on cases of Rdatasets (left: a10_ts, right: ausbeer_ts).
IV-A Noisy Almost Periodic Functions
We construct noisy almost-periodic functions by superimposing 2 trigonometric waves with incommensurate frequencies and adding Gaussian noise:
$f_{1}\left(t\right)=\cos\left(2t\right)+\cos\left(2\pi t\right)+\text{noise},$ (18)
$f_{2}\left(t\right)=\sin\left(3t\right)+\sin\left(2et\right)+\text{noise},$ (19)
where the noise is sampled from a zero-mean Gaussian distribution with variance $\sigma^{2}$. Almost-periodic functions like this are of particular significance in the development of harmonic analysis, and they form the basis of generalized harmonic analysis (GHA) as formulated by Wiener[2].
We vary the noise level $\sigma$ from 0.1 to 0.4 and compare the performance of ARIMA and 9 neural models. The results are shown in TABLE I. Typically, the outcomes of certain experiments ($\sigma=0.4$) produced by ARIMA, AR-KAN, and FAN are shown in Fig. 3.
Experimental results show that for almost-periodic functions, all 7 existing neural networks perform worse than ARIMA, including FNO and FAN, both of which are designed specifically for spectral learning. As illustrated in Fig. 3, FAN is only able to capture the rough trend of the signal but fails to reconstruct fine-grained details. In contrast, the AR-KAN achieves excellent performance comparable to ARIMA. It inherits the strong spectral analysis capabilities of autoregressive models while also benefiting from the KAN’s near absence of spectral bias, enabling it to handle the intricate details of the time series effectively.
This combination of strengths makes AR-KAN particularly suitable for data with complex frequency structures. The results highlight the effectiveness of our architecture in bridging the gap between traditional statistical methods and modern neural networks.
IV-B Rdatasets
To further demonstrate the generalizability of AR-KAN, we select 18 time series from the Rdatasets for evaluation, including a10_ts, airpass_ts, and others. These time series span a wide range of domains and exhibit diverse temporal characteristics. The performance results across all models are summarized in Table II. Typically, the outcomes of certain series (a10_ts and ausbeer_ts) produced by ARIMA, AR-KAN, and FNO are shown in Fig. 4.
The results clearly demonstrate the superior generalizability of AR-KAN across a diverse set of real-world time series. AR-KAN achieves the best performance on 13 out of 18 datasets (72%), demonstrating a notable superiority compared to the other baseline models. This consistent advantage across datasets of varying domain, scale, and statistical properties underscores AR-KAN’s capacity to adapt effectively to a wide range of temporal dynamics.
In contrast, ARIMA, despite its strong performance on the synthetic almost-periodic task, performs best on only a single real-world dataset. This sharp performance drop highlights ARIMA’s limited expressiveness when faced with the complexity and noise inherent in practical scenarios. Similarly, Fourier-based models such as FAN and FNO, though designed to leverage spectral structures, fail to generalize effectively. Their reliance on strong spectral priors appears insufficient for capturing the heterogeneous and often non-stationary nature of real-world signals.
Other deep learning models such as MLP, LSTM, and even KAN alone exhibit fluctuating performance and lack robustness across datasets. These models either overfit to local patterns or underperform due to inadequate inductive biases. By the way, with the rapid advancement of large language models (LLMs), the performance of LLMTime[42] has improved from being consistently inferior to ARIMA[41] to roughly on par with it. However, it still lags significantly behind AR-KAN. LLMs still have a long way to go in the domain of time series modeling.
By contrast, AR-KAN offers a principled blend of classical signal modeling and modern neural adaptation. It not only inherits the frequency-awareness of traditional methods, but also leverages the expressive power of neural architectures to model nonlinear and non-stationary behaviors. This versatility positions AR-KAN as a highly effective solution for a wide spectrum of time series modeling tasks.
V Conclusion
In this paper, we identified the limitations of existing neural networks in spectral analysis and demonstrated through experiments that they underperform traditional ARIMA models when dealing with almost-periodic functions. Motivated by this observation, and based in the Universal Myopic Mapping Theorem, we proposed AR-KAN, which is a novel model that effectively integrates the autoregressive component of ARIMA with the expressive nonlinear modeling capability of KAN. Experimental results show that AR-KAN achieves performance comparable to ARIMA on almost-periodic functions, and outperforms all baselines on 13 out of 18 real-world datasets, demonstrating both its effectiveness and strong generalizability.
More broadly, the contribution of this paper goes beyond AR-KAN as a single model; it brings renewed attention to Universal Myopic Mapping Theorem. This framework enables the integration of classical signal processing techniques with modern neural architectures, thereby offering interpretability and adaptability, and improving the versatility of neural models in time series tasks. We hope that future research will further explore this direction and extend the framework to more complex and diverse time series analysis scenarios.
Appendix A Data Sampling and Evaluation Protocol
In the Noisy Almost Periodic Functions experiment, the temporal variable $t$ ranges from 0 to $8\pi$, and a total of 500 samples are uniformly collected over this interval. The dataset is split into training and testing sets with an 80/20 ratio: the first 80% of the sequence is used for training, while the remaining 20% is reserved for testing.
For the Rdatasets experiment, all time series are standardized based on their mean and standard deviation. Then also apply the 80/20 split strategy: the training set consists of the first 80% of each sequence, and the testing set consists of the final 20%.
Appendix B Model Architecture and Configuration
models architecture and configuration
ARIMA p = 20, d = 0 or 1, q = 1 or 2
KAN width = [20,50,1], grid=3, k=3
MLP width = [20, 128, 256, 128, 1]
Transformer
feature_dimension = 64, nhead=4,
encoder_layers = 2, feedforward_dimension = 128
LSTM
input_size=1, hidden_size=64,
num_layers=2, output_size=1
Mamba
input_dim=1, d_model=48, d_state=32,
d_conv=20, n_layers=5
FAN
input_dim=20, output_dim=1, hidden_dim=2048,
num_layers=5, p_ratio=0.25
FNO
input_dim=20, output_dim=1, modes=8,
channels=32, fourier_layers = 2
LLMTime DeepSeek-V3, experiment_times = 10

### v2

@@
-AR-KAN: Autoregressive-Weight-Enhanced Kolmogorov–Arnold Network for Time Series Forecasting
+AR-KAN: Autoregressive-Weight-Enhanced Kolmogorov-Arnold Network for Time Series Forecasting
 Abstract
-[…]
+Traditional neural networks struggle to capture the spectral structure of complex signals.
+Fourier neural networks (FNNs) attempt to address this by embedding Fourier series components, yet many real-world signals are almost-periodic with non-commensurate frequencies, posing additional challenges.
+Building on prior work[41] showing that ARIMA outperforms large language models (LLMs) for forecasting, we extend the comparison to neural predictors and find ARIMA still superior.
+We therefore propose the Autoregressive-Weight-Enhanced Kolmogorov–Arnold Network (AR-KAN), which integrates a pre-trained AR module for temporal memory with a KAN for nonlinear representation.
+The AR module preserves essential temporal features while reducing redundancy.
+Experiments demonstrate that AR-KAN matches ARIMA on almost-periodic functions and achieves the best results on 72% of Rdatasets series, with a clear advantage on data with periodic structure.
+These results highlight AR-KAN as a robust and effective framework for time series forecasting.
+Our code is available at https://github.com/ChenZeng001/AR-KAN.
 Index Terms:
@@
 This property allows the model to flexibly adapt to various temporal patterns without manual intervention.
-Our experiments confirm the effectiveness and generalizability of AR-KAN.
-On almost-periodic functions, it matches the performance of ARIMA.
-Moreover, across 18 real-world datasets from diverse domains, AR-KAN outperforms all baselines and achieves the best results on 72% of them (13 out of 18).
-These results highlight AR-KAN’s robustness, adaptability, and promise as a unified framework for time series forecasting.
+Extensive experiments demonstrate AR-KAN’s effectiveness.
+On almost-periodic functions, it matches ARIMA’s performance.
+And on real-world datasets from Rdatasets[40], it outperforms baselines on 72% of tasks.
+The few cases where AR-KAN does not dominate correspond mostly to datasets with weak or nearly absent periodicity, which are inherently difficult to forecast without additional prior information.
+When focusing on datasets with stronger periodic components, AR-KAN shows a clear and consistent advantage, highlighting its robustness, adaptability, and potential as a unified framework for time series forecasting.
 The structure of this paper is organized as follows:
@@
 Section III presents the Universal Myopic Mapping Theorem and explains how it inspires the overall architecture of our AR-KAN model.
-Section IV describes the experiments conducted to demonstrate the effectiveness and generalizability of AR-KAN, including evaluations on two constructed almost-periodic functions and 18 real-world time series.
-Finally, Section V concludes the paper and discusses potential directions for future research.
+Section IV describes the experiments conducted to demonstrate the effectiveness and generalizability of AR-KAN, including evaluations on two constructed almost-periodic functions and real-world time series.
+Finally, Section V concludes the paper.
 II Background
@@
 First, we perform experiments on noisy almost-periodic functions to show that modern models fall short of traditional ARIMA models in terms of spectral analysis, while our AR-KAN achieves performance comparable to ARIMA.
-Then, we extend the evaluation to 18 real-world datasets from Rdatasets[40], demonstrating that AR-KAN achieves the best performance on 72% of them.
-The detailed experimental settings are provided in appendices.
-TABLE II: Test loss (MSE) of various models on Rdatasets
-Datasets ARIMA AR-KAN AR-MLP KAN MLP LSTM FAN FNO LLMTime
-a10_ts 0.1441 0.1353 0.4775 2.5033 2.2638 0.8809 0.4913 0.3851 0.3457
-airpass_ts 0.3329 0.0706 0.0871 0.3046 0.3025 0.4249 0.5163 0.6982 0.1937
-ausbeer_ts 0.0418 0.0357 0.0741 0.1031 0.5102 0.0692 0.0802 0.1114 0.0436
-auscafe_ts 0.3301 0.3813 0.1180 2.6312 0.7564 0.3746 1.4820 3.4769 0.4463
-BJsales_ts 0.3241 0.0032 0.0261 0.0358 0.7849 0.0643 0.2393 1.0370 0.0131
-bricksq_ts 0.2080 0.0502 0.0823 0.2542 0.2769 0.2961 0.9625 0.2607 0.2541
-co2_ts 0.0218 0.0014 0.0064 0.3079 0.0460 0.1640 0.1584 0.1963 0.0109
-discoveries_ts 1.6030 2.1695 2.3091 1.7269 1.6264 1.7949 1.0153 0.8469 1.2922
-economics_df_ts 3.5659 0.0845 0.4398 2.3047 1.8717 2.2520 1.8607 7.7670 1.6490
-elec_ts 0.2731 0.0069 0.0060 0.093 0.0436 0.1833 0.3258 0.0625 0.0727
-elecdaily_mts 0.4331 0.2123 0.2573 0.441 0.5366 0.6792 0.8981 0.5919 0.6127
-elecequip_ts 0.3159 0.1528 0.1346 0.5968 0.5538 0.8984 0.7761 0.4870 1.4010
-euretail_ts 0.4967 0.9964 1.3328 1.1984 0.4226 1.6740 1.1469 0.1821 1.5009
-goog200_ts 4.7135 0.1228 0.8096 3.6632 3.5888 3.2584 3.0580 7.9012 1.1351
-gtemp_both_ts 2.2374 0.2936 0.5328 3.2225 2.2946 1.6629 2.8678 1.6660 5.0938
-h02_ts 0.2726 0.1263 0.1782 1.3708 0.5258 0.2103 0.6003 0.8209 0.1371
-hsales2_ts 0.5781 0.5232 0.6301 2.1212 0.8286 1.7787 1.7065 0.8850 0.5667
-hyndsight_ts 0.8729 0.2471 0.3961 1.5734 0.4892 0.5929 0.6164 0.6793 0.4510
-Note: Bold numbers indicate the minimum value in each row.
-Figure 4: Performance of ARIMA, AR-KAN and FNO on cases of Rdatasets (left: a10_ts, right: ausbeer_ts).
+Then, we extend the evaluation to real-world datasets from Rdatasets[40], and observe that AR-KAN achieves the best performance on 72% of datasets.
+The few cases where it underperforms mostly involve weak or absent periodicity, which are inherently hard to predict without prior knowledge.
+For strongly periodic data, AR-KAN shows a clear and consistent advantage.
+TABLE II: Test loss (MSE) of various models on Rdatasets with periodicity strength
+Datasets Periodicity Strength ARIMA AR-KAN AR-MLP KAN MLP LSTM FAN FNO LLMTime
+airpass_ts 41.28% 0.3329 0.0706 0.0871 0.3046 0.3025 0.4249 0.5163 0.6982 0.1937
+goog200_ts 34.24% 4.7135 0.1228 0.8096 3.6632 3.5888 3.2584 3.0580 7.9012 1.1351
+euretail_ts 18.39% 0.4967 0.9964 1.3328 1.1984 0.4226 1.6740 1.1469 0.1821 1.5009
+ausbeer_ts 14.63% 0.0418 0.0357 0.0741 0.1031 0.5102 0.0692 0.0802 0.1114 0.0436
+h02_ts 13.31% 0.2726 0.1263 0.1782 1.3708 0.5258 0.2103 0.6003 0.8209 0.1371
+BJsales_ts 10.08% 0.3241 0.0032 0.0261 0.0358 0.7849 0.0643 0.2393 1.0370 0.0131
+a10_ts 6.36% 0.1441 0.1353 0.4775 2.5033 2.2638 0.8809 0.4913 0.3851 0.3457
+hsales2_ts 4.20% 0.5781 0.5232 0.6301 2.1212 0.8286 1.7787 1.7065 0.8850 0.5667
+co2_ts 3.69% 0.0218 0.0014 0.0064 0.3079 0.0460 0.1640 0.1584 0.1963 0.0109
+hyndsight_ts 2.61% 0.8729 0.2471 0.3961 1.5734 0.4892 0.5929 0.6164 0.6793 0.4510
+bricksq_ts 1.33% 0.2080 0.0502 0.0823 0.2542 0.2769 0.2961 0.9625 0.2607 0.2541
+elecequip_ts 1.25% 0.3159 0.1528 0.1346 0.5968 0.5538 0.8984 0.7761 0.4870 1.4010
+elecdaily_mts 1.16% 0.4331 0.2123 0.2573 0.4410 0.5366 0.6792 0.8981 0.5919 0.6127
+gtemp_both_ts 0.50% 2.2374 0.2936 0.5328 3.2225 2.2946 1.6629 2.8678 1.6660 5.0938
+discoveries_ts 0.22% 1.6030 2.1695 2.3091 1.7269 1.6264 1.7949 1.0153 0.8469 1.2922
+elec_ts 0.16% 0.2731 0.0069 0.0060 0.0930 0.0436 0.1833 0.3258 0.0625 0.0727
+economics_df_ts 0.00% 3.5659 0.0845 0.4398 2.3047 1.8717 2.2520 1.8607 7.7670 1.6490
+auscafe_ts 0.00% 0.3301 0.3813 0.1180 2.6312 0.7564 0.3746 1.4820 3.4769 0.4463
+Note: Bold numbers indicate the minimum value in each row. Rows are shaded in red when periodicity is present ($\geq 0.5\%$) and in blue when periodicity is weak ($0.5\%$).
+Figure 4: Performance of ARIMA and AR-KAN on two different types of time series in Rdatasets.
+the left column shows results on a10_ts (a highly periodic series), and the right column shows results on discoveries_ts (a nearly non-periodic series).
 IV-A Noisy Almost Periodic Functions
@@
 IV-B Rdatasets
-To further demonstrate the generalizability of AR-KAN, we select 18 time series from the Rdatasets for evaluation, including a10_ts, airpass_ts, and others.
-These time series span a wide range of domains and exhibit diverse temporal characteristics.
-The performance results across all models are summarized in Table II.
-Typically, the outcomes of certain series (a10_ts and ausbeer_ts) produced by ARIMA, AR-KAN, and FNO are shown in Fig.
+We further evaluate AR-KAN on real-world series from Rdatasets[40].
+For each dataset, we quantify its Periodicity Strength as the ratio of the energy of the seasonal component (obtained by STL decomposition[43]) to the total energy of the original series:
+$\text{Periodicity Strength}=\frac{\|x_{\text{seasonal}}\|_{2}^{2}}{\|x\|_{2}^{2}}.$
+The period used in STL decomposition is determined by the largest nonzero-lag peak in the series’ autocorrelation function.
+Our results in TABLE II show that AR-KAN achieves the best performance on 72% of datasets.
+Moreover, its advantage grows with increasing Periodicity Strength, indicating that AR-KAN excels on series with clear or strong periodic components.
+In contrast, its few underperforming cases correspond to datasets with weak or nearly absent periodicity, which are inherently hard to forecast without additional prior knowledge and remain challenging for all competing methods.
+They are illustrated in Fig.
 4.
-The results clearly demonstrate the superior generalizability of AR-KAN across a diverse set of real-world time series.
-AR-KAN achieves the best performance on 13 out of 18 datasets (72%), demonstrating a notable superiority compared to the other baseline models.
-This consistent advantage across datasets of varying domain, scale, and statistical properties underscores AR-KAN’s capacity to adapt effectively to a wide range of temporal dynamics.
-In contrast, ARIMA, despite its strong performance on the synthetic almost-periodic task, performs best on only a single real-world dataset.
-This sharp performance drop highlights ARIMA’s limited expressiveness when faced with the complexity and noise inherent in practical scenarios.
-Similarly, Fourier-based models such as FAN and FNO, though designed to leverage spectral structures, fail to generalize effectively.
-Their reliance on strong spectral priors appears insufficient for capturing the heterogeneous and often non-stationary nature of real-world signals.
-Other deep learning models such as MLP, LSTM, and even KAN alone exhibit fluctuating performance and lack robustness across datasets.
-These models either overfit to local patterns or underperform due to inadequate inductive biases.
-By the way, with the rapid advancement of large language models (LLMs), the performance of LLMTime[42] has improved from being consistently inferior to ARIMA[41] to roughly on par with it.
-However, it still lags significantly behind AR-KAN.
-LLMs still have a long way to go in the domain of time series modeling.
-By contrast, AR-KAN offers a principled blend of classical signal modeling and modern neural adaptation.
-It not only inherits the frequency-awareness of traditional methods, but also leverages the expressive power of neural architectures to model nonlinear and non-stationary behaviors.
-This versatility positions AR-KAN as a highly effective solution for a wide spectrum of time series modeling tasks.
+By the way, Fourier-based models fail to generalize despite their spectral priors, and other neural baselines show similar limitations.
+LLM-based methods such as LLMTime[42] approach ARIMA[41] in performance but still fall short of AR-KAN, indicating that LLMs are not yet mature for time series forecasting.
+Both LLMs and neural networks struggle to capture frequency structure, whereas AR offers a statistical view for genuine spectral analysis.
+By combining this frequency-aware modeling with KAN’s nonlinear expressivity, AR-KAN achieves robust, domain-general performance.
 V Conclusion
-In this paper, we identified the limitations of existing neural networks in spectral analysis and demonstrated through experiments that they underperform traditional ARIMA models when dealing with almost-periodic functions.
-Motivated by this observation, and based in the Universal Myopic Mapping Theorem, we proposed AR-KAN, which is a novel model that effectively integrates the autoregressive component of ARIMA with the expressive nonlinear modeling capability of KAN.
-Experimental results show that AR-KAN achieves performance comparable to ARIMA on almost-periodic functions, and outperforms all baselines on 13 out of 18 real-world datasets, demonstrating both its effectiveness and strong generalizability.
-More broadly, the contribution of this paper goes beyond AR-KAN as a single model; it brings renewed attention to Universal Myopic Mapping Theorem.
-This framework enables the integration of classical signal processing techniques with modern neural architectures, thereby offering interpretability and adaptability, and improving the versatility of neural models in time series tasks.
-We hope that future research will further explore this direction and extend the framework to more complex and diverse time series analysis scenarios.
+In this paper, we reveal that existing neural networks struggle with spectral analysis and often underperform ARIMA on almost-periodic functions.
+Guided by the Universal Myopic Mapping Theorem, we propose AR-KAN, which combines ARIMA’s autoregressive memory with KAN’s nonlinear expressivity.
+Experiments show that AR-KAN matches ARIMA on almost-periodic functions and outperforms baselines on 72% of real-world datasets, with its advantage growing on series with clear periodic patterns.
+These results highlight AR-KAN as a robust and unified framework for time series forecasting.
 Appendix A Data Sampling and Evaluation Protocol